OpenStudy (surana):

Calculate the heat required to melt 36.2 g of solid methanol at it's melting point.

OpenStudy (surana):

The molar mass of methanol is 32.04g. The molar enthalpy of vaporation of methanol is 35.2 kj/mol. The molar enthalpy of fusion of methanol is 3.22 kj/mol.

OpenStudy (frostbite):

If you assume that the vaporization and melting happens at the same temperature then all you got to do is to add each contribute from the vaporization and fusion together adjusted with the amount of moles of methanol in your system. Mathematically: \[\large Q= \Delta H_{fus} \times n(MetOH)+\Delta H_{vap} \times n(MetOH)\] Calculate \(n(MetOH)\) using the familiar relation \(n=m/M\)

OpenStudy (jfraser):

@frostbite vaporization and melting do \(not\) occur at the same temperature, but melting and \(fusion\) do. All you need is to apply the molar ratio of kJ/mol to the mass of methanol you have present

OpenStudy (frostbite):

\(\large Q= \Delta H_{fus} \times n(MetOH)\) I was reading incorrectly, but it still doesn't make what I say wrong! (that is the question asked us to find out how much energy is required to find the methanol in the gasphase). My approximation simply build on idea as shown below.  It is a very poor approximation, but otherwise we need to solve \[\large Q=\Delta H_{fus} \times n(MetOH) + \int\limits_{T_1}^{T_2}C_p + \Delta H_{vap} \times n(MetOH)\]  To which we don't have the information required.

OpenStudy (joannablackwelder):

Yes, what @JFraser said. Convert the mass to moles. Then use the enthalpy of fusion to convert to kJ.

OpenStudy (jfraser):

the methanol is already at its melting point, there is no need to apply a change in temperature curve. IF the methanol were being heated from solid to liquid to gas, then we'd apply the integral to find the energy required, but it's not that complicated

OpenStudy (frostbite):

I know, but as I said due to my misreading, I thought we wanted to end in gasphase, with the heat capacity not mentioned, the only way to look at it can be to: 1) assume the changes happens at the same temperature 2) assume the heat capacity is equal to 0.

OpenStudy (abb0t):

Why are you making this much more difficult than it needs to be for them? This is a high school question.

OpenStudy (surana):

I take it this is going to be a while. And I'm a he, by the way.

OpenStudy (anonymous):

I like your head dress.

OpenStudy (surana):

Thank you.

OpenStudy (surana):

Closed.

OpenStudy (surana):

Because now irrelevant.